LHC 2011 - latest news
Week 20: Machine coordinators: Gianluigi Arduini, Bernhard Holzer.
Goals of the week: re-establish physics conditions, alignment
of RPs, increase intensity
Link to report from morning meeting
- 08:00 restart, set up injection
- 07:00 Beam dump: end of STABLE BEAMS #1799: 25 pb-1.
Record integrated luminosity in 24 hours: 33.33 pb-1
- Power glitch at 09:09. Trip of sector 67 and 81
- 11:00 - 13:00. Injection. Observation of blow-up dump
and refill with lowr chormaticity
- 14:24 Stable beams #1800: 912 bunches/train - Initial
luminosity 9.4x1032 cm-2s-1
- 15:40. ALICE solenoid trip and RF dump. End of fill
#1800. 4.4 pb-1 in 1.4 h.
- 20:40 Stable beams #1801: 912 bunches/train - Initial
luminosity7.7x1032 cm-2s-1. Blow-up
observed during the ramp B1V.
- 21:27. RF dump. End of fill #1801.
- 02:13. Stable beams #1802. 912 bunches/train -
Initial luminosity1.1x1033 cm-2s-1
- 05:25. RF dump. End of stable beams #1802. 11.3
pb-1 in 2 h.
Saturday, 21st May
- 3:08 stable beams, 768 x 768. 108
b/train. Initial luminosity: 7.5x1032 cm-2s-1.
- 12:00 End of stable beams #1798. RF end
of fill test. 18.9 pb-1 in 9 hours
- 13:55 Beam dumped after RF test (switch
- 15:00-18:00 Access for ATLAS and QPS
piquet (reset of QPS controller)
- 20:34 STABLE BEAMS #1799 - 912
bunches/train - Initial
luminosity 8.5x1032 cm-2s-1.
More than 1014
p/beam in collision.
- 7:44 beam dump due to cryo problem
in point 81 (controller of DFBAA cryo-valves)
- Luminosity until then 18.5 pb-1 in 8.1 h.
- Access for replacement of the controller
for the DFBAA cryo-valves
- 12:00h restart: check injection losses
for 72 bunch transfer / 108 bunch transfer
- 14:00 injection of 768 x 768 bunches,
dump for IP6 BPM interlock test
- 15:00h injection of 768 x 768 bunches
(108 bunch trains) for
luminosity ramp,squeeze, collisions
- 16:00 stable luminosity
- 16:46 beam dump, QPS problem in
- 17:00-20:30 Investigations on Q10.L1QPS
- 20:30-23:30 trip of triplet power
Thursday, 19th May
- 00:00 - 01:00 loss maps at injection
- 01:00 - 06:00 Attempts to fill for
physics with 768 bunches/beam (72 bunches/train). Vacuum
instabilities jn the triplet L8 related to beam screen
- 06:00 - 11:00 Investigations from vacuum
- 11:20 Trip of sector 67 (QPS)
- 11:20 - 15:00 Access for QPS intervention
- 15:00-17:00 Setting-up of injection with
- 17:00-20:00 Injection of large number of
- 20:00 1308 bunch/beam injected
- 21:20 Beam dump due to vacuum spike close
- 23:30 Stable beams #1795. Peak luminosity
Wednesday, 18th May
- 01:00-06:30 Fill #1787 – 480 bunches with
72 bunches/train – 4.4x1032 cm-2s-1
initial luminosity – 3.4 pb-1 in 2.3 hours. End
of fill test of measurement of transient beam loading when
switching off a cavity.
- 06:30 - 21:30 ALPHA/TOTEM roman pot beam
- Loss maps done (except off-momentum)
- 21:30 - 24:00 - Ramp-down and
modification of the acceleration rate for the RF frequency
Tuesday, 17th May
- 01:30-03:30 Access for replacing the
power trigger unit of MKB B1
- 03:30-05:00 pre-cycle and prepare for
- 05:00-06:00 Injection of trains of 72
bunches for preparing the fill with 480 bunches. Losses at
- 06:00 Start preparation for roman port
alignment (ALPHA and TOTEM)
- 10:00 3.5 TeV - colliding. Start roman
pot alignment ...
- 10:54 beam dump due to roman pot movement
- 11:50 sector 81 & 12 tripped during
(SPS main quadrupole power supply tests)
- 13:06 end of pre-cycle, trip of sector 78
- 13:00 - 18:00 Investigations on the trips
of Sector 78 main quadrupoles.
- 18:30 finally successful pre-cycle after
restoring threshold on nQPS (B11.L8)
- 18:30 No injection permit from CMS
- 19:00 Injection of B2. Power converter
problem on RBIH.610337
- 19:30-01:00 Injection setting-up of up to
- Overnight Sunday to Monday: 50 ns inojection optimization, issues with transverse damper settings, ramp lost to electrical network glitch early morning
- 11:26 STABLE BEAMS #1785 (228 bunches/beam - 36
bunches/injection) - peak lumi 1.9x1032 cm-2s-1
- Length scale calibration in ATLAS
- 15:00 ADJUST MODE: end of fill #1785 (1.8 pb-1)
test to see effect of a trip of a klystron. Issues with logging of abort
- 17:00 Beams dumped by operator.
- 17:30 trip of triplets power converters in point 8.
Access for Power Converter piquest and for QPS piquet for replacement of
QH power supply switch in Sector 56
- 21:30 end of access.
- 22:20 Accidental CMS operator dump
- 22:30-23:30. Abort Gap Keeper check. Last bunch that
can be systematically injected.
- 23:30-01:30 Loss maps: Horizontal for B1 and B2.
Vertical for B2. Off momentum@+/-1kHz for B2.
- LBDS self-trigger. Access
VdM run summary